The base of an isosceles triangle is 14 in. less than the length of the equal sides. If the perimeter of the triangle is 43 in., find the length of each of the sides.
Hint: An isosceles triangle has 2 sides of equal length. The third side is different.
Hint: To find the perimeter of a triangle, add all 3 sides.
Hint: Be sure to check your answer. Do all three sides add up?
Answer: The length of the three sides of the triangle are ________ inches, ________ inches, and __________ inches.

Show work:

Answer by ptaylor(2198)

Let x=length of the equal sides
Then (x-14)=base
Now we are told that:
x+x+(x-14)=43 get rid of parens
x+x+x-14=43 subtract 14 from both sides
x+x+x=43+14 collect like terms
3x=57 divide both sides by 3
x=19 in---------------------------------length of the equal sides
x-14=19-14=5 in ----------------------------------length of base
CK
perimeter=43 in
19+19+5=43
43=43
also
base is 14 in less than the length of the equal sides
19-5=14
14=14
Hope this helps----ptaylor